A car is moving with 30m/s on a circular path of radius 500m .it speed is increasing at the rate of 2 metre per second square what is the acceleration of car at the moment

Grade:12th

1 Answers

hitesh
6 years ago
Given v=30a(tent. acc.)=2m/se2a(c)=?a(c)=V*2/R =30×30/500 =9/5 m/sec*2 =1.8a(net)=(2×2+1.8×1.8)^2 =7.24^2 =2.7m/sec^2